OpenStudy (anonymous):

45 is what percent of 15?

OpenStudy (lgbasallote):

try \[\frac{45}{15} \times 100\] does that help?

OpenStudy (anonymous):

yes thnxs

Parth (parthkohli):

If you are all right with some letters and numbers: Let's say that it is \(x\) percent.\[45 = 15\times {x \over 100} \]

OpenStudy (anonymous):

300%

OpenStudy (anonymous):

thanks ^

Parth (parthkohli):

You can also do it like this: 45 is three times 15. So, \(45 = {3 \over 1} \times 15\) Express \(3 \over 1\) with denominator 100.

OpenStudy (anonymous):

kk thanks